# Understanding Token Probability Encoding in Output Embeddings

### v1

@@ 7 Limitations
Although we declare that the probability editing algorithm proposed in §3 is only an experimental method for investigation, we acknowledge that it is elegant but not practical. It can never be faster, more accurate, and more harmless than a filter on the output head Guo et al. (2017). Future works can be focused on a local or directional probability editing method, limiting the detecting dataset $\mathcal{D}$, and only editing the probability on specific input prefix.
@@ Acknowledgements
This work is supported by JSPS KAKENHI Grant Number 19K20332.
The authors would like to thank Mr. Yuxuan Wang for his proofreading and constructive criticism.
Appendix A Related Works
Figure 8: The expected probability scales and the actually edited scales measured in the edited LMs w.r.t. different values of $b$.
@@ Appendix A Related Works
Figure 10: Supplement for Fig. 4. The expected probability scales and the actually edited scales measured in the edited GPT2-XL.
@@
Geometry of Input Embedding.
As a similar research object with the output embedding, it was found that the word embeddings in LMs, as well as the hidden states, are anisotropy Mu and Viswanath (2018); Ethayarajh (2019); Gao et al. (2018), i.e., these vectors share a common bias and a close direction. Such anisotropies hurt the expressiveness of word embeddings, and the word frequency in the corpus may be an inducement Mu and Viswanath (2018); Valentini et al. (2023). Also, some efforts tried to remove the harmfulness of anisotropies and towards isotropy word embeddings Mu and Viswanath (2018); Gong et al. (2018). These works are based on input embedding, while our paper is on output embedding. Although we can confirm that the input and output embeddings act similarly, they are still completely different components of untied LMs. So the existing conclusions on input embeddings cannot overwrite our work.
Embedding Tying in LMs.
(Language) Model Editing and Model Pruning.
Training Dynamics (of LMs).
Investigating what is happening in the training process of language models and other deep learning models is an attractive research topic. Many works about training trajectory Kalra and Barkeshli (2024); Jastrzębski et al. (2018); Lewkowycz et al. (2020), early period training behaviors Frankle et al. (2019); Kalra and Barkeshli (2024); Achille et al. (2018), loss landscape Neyshabur et al. (2020); Keskar et al. (2017); Li et al. (2018), and "knowledge" earned in different stages of training Tirumala et al. (2022); Liu et al. (2021) have been done. Differences in training speed among various network layers (the deeper-slower pattern) have been discovered by Chen et al. (2022), while in this paper, following their method, we don’t find a similar pattern.
@@
(a) $e_{local}$
(b) $e_{id}$
(c) $e_{ood}$
(d) $d_{KL}^{r}$
@@
Figure 13: Fine-grained results w.r.t. expected editing scales of GPT2.
(a) $e_{local}$
(b) $e_{id}$
(c) $e_{ood}$
(d) $d_{KL}^{r}$
@@
Figure 14: Fine-grained results w.r.t. expected editing scales of GPT2-XL.
(a) $e_{local}$
(b) $e_{id}$
(c) $e_{ood}$
(d) $d_{KL}^{r}$
Figure 15: Fine-grained results w.r.t. expected editing scales of GPT-J.
(a) $e_{local}$
(b) $e_{id}$
(c) $d_{KL}^{r}$
Figure 16: Supplement for Fig. 5. The $e_{local}$, $e_{id}$, $d_{KL}^{r}$ w.r.t. the detect dataset size.
@@
Appendix B Calculation Details
B.1 Calculation of Averaged Token Probability
B.2 Error Analysis of Eq. 1
In Eq. 1, we do a local linear approximation as:
$\begin{split}-\log\alpha_{w,\mathcal{D},\theta}&=-\log\mathbb{E}_{x\in\mathcal{D}}[P_{\theta}(w|x)]\\ &\approx\mathbb{E}_{x\in\mathcal{D}}[-\log[P_{\theta}(w|x)]].\\ \end{split}$
That is, given a set of $\{p_{i}\}_{i=0}^{n}$, where $\forall i,p_{i}>0$ we approximate that $\log\mathbb{E}_{i\in[0,n]}[p_{i}]\approx\mathbb{E}_{i\in[0,n]}[\log p_{i}]$. Assume that we have a non-descending sequence of $p$, that is, $p_{0}\leq p_{1}\leq p_{2}\leq\dots\leq p_{n}$, w.l.o.g. We can confirm that $\mathbb{E}_{i\in[0,n]}[p_{i}]\in[p_{0},p_{n}]$. So we have: $\exists\xi\in(p_{0},\mathbb{E}_{i\in[0,n]}[p_{i}]),s.t.$,
We have:
$\mathbb{E}_{i\in[0,n]}[\log p_{i}]\geq\log p_{0},$
since $\log^{\prime}(\cdot)>0$. That is:
$\begin{split}&[\log\mathbb{E}_{i\in[0,n]}[p_{i}]-\mathbb{E}_{i\in[0,n]}[\log p_{i}]]^{2}\\ \leq&\frac{1}{\xi^{2}}(\mathbb{E}_{i\in[0,n]}[p_{i}]-p_{0})^{2}\\ \leq&\frac{1}{p_{0}^{2}}(\mathbb{E}_{i\in[0,n]}[p_{i}]-p_{0})^{2}.\end{split}$ (3)
B.3 Details of MAUVE Calculation
The MAUVE is a similarity between two datasets. Refer Pillutla et al. (2021) for the calculating details of MAUVE. Here we explain our method to get both datasets for MAUVE to measure the harmfulness of our model editing method. We sample 512 data points from the BookCorpus Zhu et al. (2015), and take the first two words of each data point as a prefix to collect a prefix set. Given a generative language model, we input the prefixes and let the model generate naturally into a generated dataset. We do this process on the original model and the edited model and use the two generated datasets to calculate the MAUVE.
Appendix C Ablation Study on Editing Amount Allocation Parameter $b$
We discuss the editing amount allocation parameter $b$ in Algorithm 1. We try different settings of $b$ on GPT2 as shown in Table 3 and Fig. 8.
The results show that when the $b$ is not infinities, the editing remains accurate and stable. That is, the algorithm is not sensitive to a positive integer $b$. But we still recommend a large $b$ when the model is large and the sparsity of MLR slope is not fine.
However, when the $b$ is +$\infty$+ ∞ or $\infty$- ∞, the editing can not be accurate and is easy to get a larger scale than expected. We infer that both of these infinity $b$ increase the norm of the embedding vector too much, leading to an increase in its output probability as well.
Appendix D Supplementary Experiment Results
@@
Supplement for Table 2. Fine-grained results w.r.t. expected editing scales of GPT2 (Fig. 13), GPT2-XL (Fig. 14), GPT-J (Fig. 15).
Supplement for Fig. 5. The $e_{local}$, $e_{id}$, $d_{KL}^{r}$ w.r.t. the detect dataset size are shown in Fig. 16.
@@
Appendix E Necessary Statements
@@
The artifacts used in this paper are all open-sourced and are used for their intended usage.
Models. GPT2 family are under the MIT license, GPT-J and Pythia are under the apache-2.0 license.
Datasets. WikiDpr is under the cc-by-nc-4.0 license, BookCorpus and Pile is under the MIT license.
Tool. MAUVE is under the GNU license.

### v2

@@ 8 Limitations
Although we declare that the probability steering algorithm proposed in §3 is only an experimental method for investigation, we acknowledge that it is elegant but not practical. It can never be faster, more accurate, and more harmless than a filter on the output head Guo et al. (2017). Future works can be focused on a local or directional probability editing method, limiting the detecting dataset $\mathcal{D}$, and only editing the probability on specific input prefix to resolve a controlled-generation task.
@@ Acknowledgements
This work is supported by the Nakajima Foundation.
The authors would like to thank Mr. Yuxuan Wang at Beijing Institute of Technology for his proofreading and constructive criticism.
Appendix A Calculation Details
A.1 Calculation of Averaged Token Probability
A.2 Error Analysis of Eq. 1
In Eq. 1, we do a local linear approximation as:
$\begin{split}-\log\alpha_{w,\mathcal{D},\theta}&=-\log\mathbb{E}_{x\in\mathcal{D}}\left[\mathrm{P}_{\theta}(w|x)\right]\\ &\approx\mathbb{E}_{x\in\mathcal{D}}\left[-\log\left[\mathrm{P}_{\theta}(w|x)\right]\right].\\ \end{split}$
That is, given a set of $\{p_{i}\}_{i=0}^{n}$, where $\forall i,p_{i}>0$ we approximate that $\log\mathbb{E}_{i\in[0,n]}[p_{i}]\approx\mathbb{E}_{i\in[0,n]}[\log p_{i}]$. Assume that we have a non-descending sequence of $p$, that is, $p_{0}\leqslant p_{1}\leqslant p_{2}\leqslant\dots\leqslant p_{n}$, w.l.o.g. We can confirm that $\mathbb{E}_{i\in[0,n]}[p_{i}]\in[p_{0},p_{n}]$. So we have: $\exists\xi\in(p_{0},\mathbb{E}_{i\in[0,n]}[p_{i}]),s.t.$,
We have:
$\mathbb{E}_{i\in[0,n]}[\log p_{i}]\geqslant\log p_{0},$
since $\log^{\prime}(\cdot)>0$. That is:
$\begin{split}&[\log\mathbb{E}_{i\in[0,n]}[p_{i}]-\mathbb{E}_{i\in[0,n]}[\log p_{i}]]^{2}\\ \leqslant&\frac{1}{\xi^{2}}\left(\mathbb{E}_{i\in[0,n]}[p_{i}]-p_{0}\right)^{2}\\ \leqslant&\frac{1}{p_{0}^{2}}\left(\mathbb{E}_{i\in[0,n]}[p_{i}]-p_{0}\right)^{2}.\end{split}$ (3)
Figure 8: The expected probability scales and the actually steered scales measured in the edited LMs w.r.t. different values of $b$.
@@ A.1 Calculation of Averaged Token Probability
Figure 10: Supplement for Fig. 5. The expected probability scales and the actually steered scales measured in the steered GPT2-XL.
@@
(a) $e_{\mathrm{local}}$
(b) $e_{\mathrm{id}}$
(c) $e_{\mathrm{ood}}$
(d) $d_{\mathrm{KL}}^{r}$
@@
Figure 13: Fine-grained results w.r.t. expected steering scales of GPT2.
(a) $e_{\mathrm{local}}$
(b) $e_{\mathrm{id}}$
(c) $e_{\mathrm{ood}}$
(d) $d_{\mathrm{KL}}^{r}$
@@
Figure 14: Fine-grained results w.r.t. expected steering scales of GPT2-XL.
(a) $e_{\mathrm{local}}$
(b) $e_{\mathrm{id}}$
(c) $e_{\mathrm{ood}}$
(d) $d_{\mathrm{KL}}^{r}$
Figure 15: Fine-grained results w.r.t. expected steering scales of GPT-J.
(a) $e_{\mathrm{local}}$
(b) $e_{\mathrm{id}}$
(c) $d_{\mathrm{KL}}^{r}$
Figure 16: Supplement for Fig. 4. The $e_{\mathrm{local}}$, $e_{\mathrm{id}}$, $d_{\mathrm{KL}}^{r}$ w.r.t. the detect dataset size.
@@
A.3 Details of MAUVE Calculation
The MAUVE is a similarity between two datasets. Refer Pillutla et al. (2021) for the calculating details of MAUVE. Here we explain our method to get both datasets for MAUVE to measure the harmfulness of our model steering method. We sample 512 data points from the BookCorpus Zhu et al. (2015), and take the first two words of each data point as a prefix to collect a prefix set. Given a generative language model, we input the prefixes and let the model generate naturally into a generated dataset. We do this process on the original model and the steered model and use the two generated datasets to calculate the MAUVE.
Appendix B Ablation Study on Steering Amount Allocation Parameter $b$
We discuss the steering amount allocation parameter $b$ in Algorithm 1. We try different settings of $b$ on GPT2 as shown in Table 3 and Fig. 8.
The results show that when the $b$ is not infinities, the steering remains accurate and stable. That is, the algorithm is not sensitive to a positive integer $b$. But we still recommend a large $b$ when the model is large and the sparsity of MLR slope is not fine.
However, when the $b$ is +$\infty$+ ∞ or $\infty$- ∞, the steering can not be accurate and is easy to get a larger scale than expected. We infer that both of these infinity $b$ increase the norm of the embedding vector too much, leading to an increase in its output probability as well.
Appendix C Supplementary Experiment Results
@@
Supplement for Table 2. Fine-grained results w.r.t. expected steering scales of GPT2 (Fig. 13), GPT2-XL (Fig. 14), GPT-J (Fig. 15).
Supplement for Fig. 4. The $e_{\mathrm{local}}$, $e_{\mathrm{id}}$, $d_{\mathrm{KL}}^{r}$ w.r.t. the detect dataset size are shown in Fig. 16.
@@
Appendix D Necessary Statements
Author Contributions.
Hakaze Cho, also pronounced Yufeng Zhao, handled virtually all the workload in this paper. N.I. is their supervisor, providing important and beneficial guidance, revision, and research support. Y.S., K.T, and M.K. participated in the discussion and provided some helpful suggestions and revisions.
@@ Appendix D Necessary Statements
The artifacts used in this paper are all open-sourced and are used for their intended usage. Models. GPT2 family are under the MIT license, GPT-J and Pythia are under the apache-2.0 license. Datasets. WikiDpr is under the cc-by-nc-4.0 license, BookCorpus and Pile is under the MIT license. Tool. MAUVE is under the GPLv3 license.